OpenStudy (anonymous):

After 8 years, what is the total amount of a compound interest investment of $25,000 at 3% interest, compounded quarterly? $25,377.11 $25,759.92 $26,148.50 $31,752.78

OpenStudy (misty1212):

hi again!

OpenStudy (anonymous):

hey lol you are just too smart haha

OpenStudy (misty1212):

you need a calculator for this one dear

OpenStudy (anonymous):

i got one but what is the formula i need to do

OpenStudy (misty1212):

\[25000\times \left(1+\frac{.03}{4}\right)^{8\times 4}\]

OpenStudy (anonymous):

misty1212 I need help in something can u help me after u help brittny

OpenStudy (misty1212):

be careful when you put it in best to know \(4\times 8=32\) first and use \[25000\times \left(1+\frac{.03}{4}\right)^{32}\]

OpenStudy (anonymous):

you are awesome thanks

OpenStudy (anonymous):

The bird population on an island is declining at a rate of 2.2% per year. If the population was 3500 in the year 2009, which is the best prediction of the population in the year 2014? 2730 3062 3132 3423 Last question then i move onto next lol

OpenStudy (misty1212):

\[\color{magenta}\heartsuit\]

OpenStudy (misty1212):

\[3500\times (.987)^5\] should do it

OpenStudy (misty1212):

@1rarity1 what do you need?

OpenStudy (anonymous):

ill tag u

OpenStudy (misty1212):

http://www.wolframalpha.com/input/?i=3500*.987^5